OpenStudy (anonymous):

The zero of the function in 9x^2-25/3x+5 is a. -5/3 b. 5/3 c. 5 **My Answer was "a" please explain how you got your answer.

OpenStudy (anonymous):

Is the problem (9x^2-25)/(3x+5)?

OpenStudy (anonymous):

Yes.

OpenStudy (whpalmer4):

Please do explain how you got that, so we can fix your thinking about this problem.

OpenStudy (whpalmer4):

were you trying to make the denominator be 0?

OpenStudy (anonymous):

There is no= in this problem, so I don't see how you can find a zero.

OpenStudy (whpalmer4):

that's called a pole, not a zero. a zero will have the numerator be equal to 0.

OpenStudy (anonymous):

by factoring 9 and 25

OpenStudy (whpalmer4):

but at x = -5/3, the function is undefined due to division by 0. that cant be your zero, you'll have to find another. how many solutions are there to 9x^2-25 = 0?

OpenStudy (anonymous):

Here ill show you how it looks on paper.

OpenStudy (anonymous):

Ah, so you are trying to find when the numerator is zero. After the factors of (3x+5) cancel, you are left with (3x-5) and 3x-5=0 when x=5/3

OpenStudy (anonymous):

Why are they positive numbers if I may ask.

OpenStudy (anonymous):

You see that the 3x+5's from the numerator and denominator will cancel, right? And this will leave us with 3x-5. Now just set 3x-5=0. Move the 5 to the other side, making it positive 5, and then divide by 3, making x=5/3.

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

You are welcome!

OpenStudy (anonymous):

Would I graph that answer?

OpenStudy (anonymous):

You can. If you have access to a graphing calculator, it will show you where you have a 'pole' as whalmer pointed out earlier at x=-5/3.

OpenStudy (anonymous):

alright.